# HSR-Enhanced Sparse Attention Acceleration

### v1

@@ 8 Discussion and Future Work
The sparsity within neural networks arises primarily from the incorporation of non-linear activation functions. These non-linear functions determine the mechanism or circuit of the neural networks, e.g., the induction head in transformers [54]. Gaining insight into these non-linear layers not only enhances our understanding of how neural networks work but also paves the way for optimizing training and inference. We hope our analysis may inspire efficient neural network architecture design. This work represents the initial point of this envisioned blueprint. We concentrate on analyzing the combinations of LLMs and fundamental non-linear activation functions—ReLU and Softmax, which are most relevant to contemporary applications. By analyzing these functions, we aim to demonstrate to the research community that a thorough examination of a model’s non-linear characteristics can significantly enhance the running time complexity of neural networks.
In real-world scenarios, a multitude of non-linear activation functions exist beyond ReLU and Softmax, such as those designated as $\mathsf{SELU}(x)=\mathrm{scale}\cdot(\max(0,x)+\min(0,\alpha\cdot(\exp(x)-1)))$ [36], $\mathsf{CELU}(x)=\max(0,x)+\min(0,\alpha\cdot(\exp(x/\alpha)-1))$ [8], and $\mathsf{PRELU}(x)=\max(0,x)+\mathrm{weight}\cdot\min(0,x)$ [30]. However, analyzing these alternative functions poses multiple challenges. Hence, we will explore these additional functions in the future.
@@ 9 Conclusion
This work investigates the exploitation of the intrinsic sparsity present in both ReLU and Softmax attention mechanisms to decrease the computational complexity of full attention computation and attention generation scenarios. Specifically, we employ the Half-Space Reporting (HSR) data structure to accelerate the process of identifying non-zero or “massive activated” entries within ReLU and Softmax attentions, respectively. Importantly, our approach does not import any errors to ReLU attention, and it results in only a negligible approximation error for Softmax attention.
@@ Acknowledgement
Roadmap. In Section A, we introduce more fundamental lemmas and facts. In Section B, we extend the analysis to ReLU attention calculation, demonstrating improved performance over standard attention computation under specific conditions. In Section C, we first introduce and analyze the time complexity of ReLU attention generation using half-space reporting (HSR) data structures. In Section D, we analyze the sparsity of ReLU attention matrices. In Section E, we introduce our results on reducing the running time of Softmax attention. In Section F, we analyze error bounds for Softmax attention with index sets, balancing efficiency and accuracy.
Appendix A Preliminary
In this section, we display more fundamental concepts. In Section A.1, we introduce several important probability properties and bounds. In Section A.2, we detail the time complexity and performance of half-space reporting (HSR) data structures.
A.1 Probability Tools
@@
Fact A.1 (Weighted summation of Gaussian).
@@
Fact A.2 (Independence between $\langle x,y_{i}\rangle$ and $\langle x,y_{j}\rangle$).
@@
Lemma A.3 (Chi-square tail bound, Lemma 1 in [43] ).
@@
Fact A.4 (Gaussian tail bound).
@@
Then we consider the numerator and we use $y=x-\mu$ to simplify the calculation, we have
@@
where the first step follows from Eq. (A.1) and Eq.(A.1), the second step follows from the calculation of infimum. ∎
@@
Lemma A.5 (Bernstein inequality [10]).
@@
A.2 Half-Space Reporting (HSR) Data Structures
@@
Theorem A.6 (Agarwal, Eppstein and Matousek [1]).
@@
Corollary A.7 (HSR data-structure time complexity [1], formal version of Corollary 3.5).
@@
Appendix B Full ReLU Attention Computation
@@
Lemma B.1 (General full attention computation framework, formal version of Lemma 6.3).
@@
  • Let hsr data structure be defined as Part 1 in Corollary A.7.
@@
The first term ${\cal T}_{\mathsf{init}}(n,d)$ corresponds to the initialization of the hsr data structure. Since we use Part 1 result from Corollary A.7, the running time for initialization is ${\cal T}_{\mathsf{init}}(m,d)=O_{d}(m\log m)$.
The second term $\sum_{i=1}^{m}{\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ comes from the HSR query operation (Line 11). Since we use Part 1 result from Corollary A.7, we have
@@
where the first step follows from ${\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})=O(dn^{1-\lfloor d/2\rfloor}+d\widetilde{k}_{i})$ (Part 1 of Corollary A.7), the second step follows from with high probability $\widetilde{k}_{i}$ at most $n^{4/5}$ (Lemma D.3).
The third term $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $A_{j,i}$ (Line 13). By Lemma D.3, we have the third term is $O(mn^{4/5})$.
The fourth term $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $D^{-1}AV$. Since for $i$-th row of $A$, there are $\widetilde{k}_{i}$ non-zero entries. Therefore, it takes $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ time for calculating $D^{-1}A$. Therefore, it takes $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$ time to calculate $D^{-1}AV$. By Lemma D.3, with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$. Therefore, we have the third term as $O(mn^{4/5}d)$.
@@
Theorem B.2 (Running time of full ReLU attention computation, formal version of Lemma 5.1).
@@
By Lemma B.1, we have that the FullAttentionComputation data structure (Algorithm 3) can run Inference to calculate the ReLU attention, in $O(m^{1-\lfloor d/2\rfloor}nd+mn^{4/5}d)$ time.
@@
Appendix C ReLU Attention Generation
In this section, we present a theoretical analysis of the time complexity of ReLU attention generation using a HSR data structure.
Lemma C.1 (General attention generation framework, formal version of Lemma 6.2).
@@ Appendix C ReLU Attention Generation
  • Let hsr data structure be defined as Part 2 in Corollary A.7.
@@
The Init procedure only runs the initialization of the HSR data structure. Since we use Part 2 result from Corollary A.7, the running time of Init procedure is ${\cal T}_{\mathsf{init}}(n,d)=O(n^{\lfloor d/2\rfloor})$.
@@
The first term $\sum_{i=1}^{m}{\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ corresponds to the HSR query operation (Line 16). Since we use the Part 2 result from Corollary A.7, we have
@@
where the first step follows from ${\cal T}_{\mathsf{query}}(n,d,k)=O(d\log n+dk)$ in Part 2 of Corollary A.7, the second step follows from with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$ (Lemma D.3), the third step follows from $\log n<n^{4/5}$.
The second term $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $A_{i,j}$ (Line 18). There are $m$ iterations, and in each iteration, it calculates $\widetilde{k}_{i}$ entries of $A$. Then, the second term is $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$. By Lemma D.3, with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$. Therefore, we have the second term as $O(mn^{4/5}d)$.
Similar to the proof of Lemma B.1 this term is $O(mn^{4/5}d)$.
@@
Theorem C.2 (Running time of full ReLU attention generation, formal version of Theorem 4.1).
@@
We make use of the AttentionGeneration data structure (Algorithm 2) in Lemma C.1.
@@
For calculating the attention between $q_{i}$ and $K\in\mathbb{R}^{n\times d}$, we just need to call AttentionGeneration .Inference$(q_{i},1)$ for once. Therefore the running time for this part is $O(n^{4/5}d)$ time.
@@
Appendix D Sparsity Analysis
@@
By the property of Gaussian distribution (Fact A.1), we have $\langle x,K_{i,*}\rangle\sim\mathcal{N}(0,d\cdot\sigma_{a}^{2})$ and $\langle x,K_{i,*}\rangle/\sqrt{d}\sim\mathcal{N}(0,\sigma_{a}^{2})$.
For any $i,j\in[n]$, by Fact A.2, we have $\langle x,K_{i,*}\rangle$ and $\langle x,K_{j,*}\rangle$ are independent, which implies $r_{i}$ and $r_{j}$ are independent.
By the tail bound of Gaussian distribution (Fact A.4), we have
@@
Let $k_{0}:=n\cdot\exp(-\frac{b^{2}}{2\sigma_{a}^{2}})$. By the Bernstein inequality (Lemma A.5), we have
@@
Since $\|x\|_{2}^{2}=\sum_{i=1}^{d}x_{i}^{2}$, by Chi-square tail bound (Lemma A.3), we have
@@
Lemma D.3 (Sparsity analysis, formal version of Lemma 6.1).
@@
This proof follows from applying union bound on Lemma D.1 and Lemma D.2.
By Lemma D.2, we have
@@
Appendix E Running Time of Softmax Attention
@@
Theorem E.1 (Running time of Softmax attention generation, formal version of Theorem 4.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 2 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under the Softmax attention generation scenario: Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $m=\Theta(1)$. Algorithm 2 (replacing ReLU attention with Softmax attention) takes $O(mn^{4/5})$ time to generate the answer.
@@
The Softmax attention generation scenario can be proved by substituting the ReLU attention $\mathsf{Attn}_{r}$ (Definition 1.2) with Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ (Definition 3.2) in Algorithm 2 and Theorem 4.1. ∎
@@
Theorem E.2 (Running time of Softmax full attention computation, formal version of Theorem 5.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 3 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under full Softmax attention computation scenario: Suppose we have $m=\Theta(n)$. Algorithm 3 (replacing ReLU attention with Softmax attention) takes $O(n^{2-1/\lfloor d/2\rfloor}d+n^{1+4/5}d)$ time to calculate the attention output.
@@
The Softmax full attention computation scenario can be proved by substituting the ReLU attention $\mathsf{Attn}_{r}$ (Definition 1.2) with Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ (Definition 3.2) in Algorithm 3 and Theorem 5.1. ∎
Appendix F Error Analysis of Softmax Attention

### v2

@@ 8 Discussion and Future Work
The sparsity within neural networks arises primarily from the incorporation of non-linear activation functions. These non-linear functions determine the mechanism or circuit of the neural networks [68]. Gaining insight into these non-linear layers not only enhances our understanding of how neural networks work but also paves the way for optimizing training and inference. We believe our analysis may inspire efficient neural network architecture design. This work represents the initial point of this envisioned blueprint. We concentrate on analyzing the combinations of LLMs and fundamental non-linear activation functions, ReLU and Softmax. By analyzing these functions, we aim to demonstrate to the research community that a thorough examination of a model’s nonlinear characteristics can significantly enhance neural networks’ running-time complexity.
In real-world scenarios, various non-linear activation functions exist beyond ReLU and Softmax, such as $\mathsf{SELU}(x)=\mathrm{scale}\cdot(\max(0,x)+\min(0,\alpha\cdot(\exp(x)-1)))$ [47], $\mathsf{CELU}(x)=\max(0,x)+\min(0,\alpha\cdot(\exp(x/\alpha)-1))$ [8], and $\mathsf{PRELU}(x)=\max(0,x)+\mathrm{weight}\cdot\min(0,x)$ [39]. Analyzing these alternative functions poses multiple challenges. We will explore these additional functions in the future.
@@ 9 Conclusion
This work investigates the exploitation of the intrinsic sparsity present in both ReLU and Softmax attention mechanisms to decrease the computational complexity of generation decoding and prompt prefilling scenarios. We employ the HSR data structure to accelerate the process of identifying non-zero or “massive activated” entries within ReLU and Softmax attentions. Our approach results in only a negligible approximation error for Softmax attention.
@@ Acknowledgement
Roadmap. In Section A, we discuss some limitations of our work. In Section B, we introduce more fundamental lemmas and facts. In Section C, we extend the analysis to ReLU attention calculation, demonstrating improved performance over standard attention computation under specific conditions. In Section D, we first introduce and analyze the time complexity of ReLU attention generation using half-space reporting (HSR) data structures. In Section E, we analyze the sparsity of ReLU attention matrices. In Section F, we introduce our results on reducing the running time of Softmax attention. In Section G, we analyze error bounds for Softmax attention with index sets, balancing efficiency and accuracy.
Sequence Activated Sparsity
length entries ratio
$1$k 251 0.75
$2$k 437 0.78
$4$k 761 0.81
$8$k 1325 0.83
$16$k 2308 0.86
$32$k 4019 0.87
$64$k 6997 0.89
$128$k 12183 0.90
$256$k 21212 0.92
$512$k 36933 0.93
$1024$k 64304 0.94
Table 1: An illustration of the sparsity level attained by our algorithm across varying sequence lengths, $n$. Our approach activates merely $n^{4/5}$ entries per inference, resulting in a computational savings of up to $90\%$ when $n=1024$k.
Appendix A Limitations
Our work is fundamentally theoretical in nature. We concentrate on the theoretical analysis of our proposed algorithms (Algorithms 1 and 2). Our study does not include an implementation of the suggested algorithms, a limitation that arises from the absence of an existing implementation for the original HSR data structure, initially proposed in [1]. We are confident that our theoretical insights will inspire the development of future algorithmic designs.
Appendix B Full Background and Definition
In this section, we display more fundamental concepts. In Section B.1, we introduce a modified version of Softmax attention that operates on a specific subset of indices. It defines the top-$r$ nearest neighbors Softmax attention, which focuses on the most relevant entries in the attention matrix. In Section B.2, we describe the massive activation property for attention mechanisms. In Section B.3, we introduce several important probability properties and bounds. In Section B.4, we detail the time complexity and performance of half-space reporting (HSR) data structures.
B.1 Softmax Attention with Index Set
Definition B.1 (Input with index set).
We consider calculating the Softmax attention on the “massive activation” index set, where we define the “massive activation” index set as the top-$r$ indices, where the formal definition is as follows:
Definition B.2 ( Top-$r$ indices Softmax attention ).
Let $q\in\mathbb{R}^{d}$, $K,V\in\mathbb{R}^{n\times d}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK$, where $|\mathsf{NN}(r,q,K)|=r$. Let $\widehat{K},\widehat{V}\in\mathbb{R}^{r\times d}$ and $\overline{K},\overline{V}\in\mathbb{R}^{(n-r)\times d}$ be defined in Definition B.1. We define the top-$r$ nearest neighbors (NN) Softmax attention computation $\widehat{\mathsf{Attn}}_{s}(q,K,V)\in\mathbb{R}^{d}$ as follows:
where
In Definition B.2, we view the “massive activated” entries as the top-$r$ entries. Therefore, we only calculate the Softmax attention based on $\widehat{K},\widehat{V}\in\mathbb{R}^{r\times d}$, instead of $K,V\in\mathbb{R}^{n\times d}$.
B.2 Massive Activation
Definition B.3 (Massive activation property).
Let $\gamma\in[0,1]$, $\beta_{1}\geq\beta_{2}\geq 0$. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK^{\top}$. We define $(\gamma,\beta_{1},\beta_{2})$ massive activation for a query $q\in\mathbb{R}^{d}$ and key cache $K\in\mathbb{R}^{n\times d}$, if the following conditions hold:
An intuitive understanding of Definition B.3 is that the summation of “massive activated” entries dominates the summation of all entries, and the entries we ignored only contribute little to the final summation. Therefore, it is reasonable for us to omit those non “massive activated” entries.
Remark B.4.
There are many distributions satisfying the property in Definition B.3, such as (1) $K$ drawing from any subexponential distribution, e.g., multivariate Laplace distributions, (2) $K$ drawing from any mixture of Gaussian distribution with $n^{1-\gamma}$ Gaussian clusters.
B.3 Probability Tools
@@
Fact B.5 (Weighted summation of Gaussian).
@@
Fact B.6 (Independence between $\langle x,y_{i}\rangle$ and $\langle x,y_{j}\rangle$).
@@
Lemma B.7 (Chi-square tail bound, Lemma 1 in [56] ).
@@
Fact B.8 (Gaussian tail bound).
@@
Then we consider the numerator, and we use $y=x-\mu$ to simplify the calculation, we have
@@
where the first step follows from Eq. (B.3) and Eq.(B.3), the second step follows from the calculation of infimum. ∎
@@
Lemma B.9 (Bernstein inequality [10]).
@@
B.4 Half-Space Reporting (HSR) Data Structures
Algorithm 3 Half Space Report Data Structure
1:data structure HalfSpaceReport
4:end data structure
We restate the result from [1] for solving the half-space range reporting problem. The half-space range reporting problem is a fundamental problem in computational geometry and can be formally defined as follows:
Definition B.10 (Half-space range reporting [1, 82]).
@@ B.4 Half-Space Reporting (HSR) Data Structures
Theorem B.11 (Agarwal, Eppstein and Matousek [1]).
@@
Corollary B.12 (HSR data-structure time complexity [1], formal version of Corollary 3.1).
@@
Appendix C ReLU Attention Prompt Prefilling
@@
Lemma C.1 (General full attention computation framework, formal version of Lemma 6.3).
@@
  • Let hsr data structure be defined as Part 1 in Corollary B.12.
@@
The first term ${\cal T}_{\mathsf{init}}(n,d)$ corresponds to the initialization of the hsr data structure. Since we use the Part 1 result from Corollary B.12, the running time for initialization is ${\cal T}_{\mathsf{init}}(n,d)=O_{d}(n\log n)$.
The second term $\sum_{i=1}^{m}{\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ comes from the HSR query operation (Line 11). Since we use Part 1 result from Corollary B.12, we have
@@
where the first step follows from ${\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})=O(dn^{1-\lfloor d/2\rfloor}+d\widetilde{k}_{i})$ (Part 1 of Corollary B.12), the second step follows from with high probability $\widetilde{k}_{i}$ at most $n^{4/5}$ (Lemma E.3).
The third term $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $A_{j,i}$ (Line 13). By Lemma E.3, we have the third term is $O(mn^{4/5})$.
The fourth term $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $D^{-1}AV$. Since for $i$-th row of $A$, there are $\widetilde{k}_{i}$ non-zero entries. Therefore, it takes $O(\sum_{i=1}^{m}\widetilde{k}_{i})$ time for calculating $D^{-1}A$. Therefore, it takes $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$ time to calculate $D^{-1}AV$. By Lemma E.3, with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$. Therefore, we have the third term as $O(mn^{4/5}d)$.
@@
Theorem C.2 (Running time of full ReLU attention computation, formal version of Lemma 5.1).
@@
By Lemma C.1, we have that the FullAttentionComputation data structure (Algorithm 2) can run Inference to calculate the ReLU attention, in $O(m^{1-\lfloor d/2\rfloor}nd+mn^{4/5}d)$ time.
@@
Appendix D ReLU Attention Generation Decoding
In this section, we present a theoretical analysis of the time complexity of ReLU attention generation using an HSR data structure.
Lemma D.1 (General attention generation framework, formal version of Lemma 6.2).
@@ Appendix D ReLU Attention Generation Decoding
  • Let hsr data structure be defined as Part 2 in Corollary B.12.
@@
The Init procedure only runs the initialization of the HSR data structure. Since we use Part 2 result from Corollary B.12, the running time of Init procedure is ${\cal T}_{\mathsf{init}}(n,d)=O(n^{\lfloor d/2\rfloor})$.
@@
The first term $\sum_{i=1}^{m}{\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ corresponds to the HSR query operation (Line 16). Since we use the Part 2 result from Corollary B.12, we have
@@
where the first step follows from ${\cal T}_{\mathsf{query}}(n,d,k)=O(d\log n+dk)$ in Part 2 of Corollary B.12, the second step follows from with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$ (Lemma E.3), the third step follows from $\log n<n^{4/5}$.
The second term $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$ corresponds to calculating $A_{i,j}$ (Line 18). There are $m$ iterations, and in each iteration, it calculates $\widetilde{k}_{i}$ entries of $A$. Then, the second term is $O(d\sum_{i=1}^{m}\widetilde{k}_{i})$. By Lemma E.3, with high probability, $\widetilde{k}_{i}$ is at most $n^{4/5}$. Therefore, we have the second term as $O(mn^{4/5}d)$.
Similar to the proof of Lemma C.1 this term is $O(mn^{4/5}d)$.
@@
Theorem D.2 (Running time of full ReLU attention generation, formal version of Theorem 4.1).
@@
We make use of the AttentionGeneration data structure (Algorithm 1) in Lemma D.1.
@@
For calculating the attention between $q_{i}$ and $K\in\mathbb{R}^{n\times d}$, we just need to call AttentionGeneration .Inference$(q_{i},1)$ for once. Therefore, the running time for this part is $O(n^{4/5}d)$ time.
@@
Appendix E Sparsity Analysis
@@
By the property of Gaussian distribution (Fact B.5), we have $\langle x,K_{i,*}\rangle\sim\mathcal{N}(0,d\cdot\sigma_{a}^{2})$ and $\langle x,K_{i,*}\rangle/\sqrt{d}\sim\mathcal{N}(0,\sigma_{a}^{2})$.
For any $i,j\in[n]$, by Fact B.6, we have $\langle x,K_{i,*}\rangle$ and $\langle x,K_{j,*}\rangle$ are independent, which implies $r_{i}$ and $r_{j}$ are independent.
By the tail bound of Gaussian distribution (Fact B.8), we have
@@
Let $k_{0}:=n\cdot\exp(-\frac{b^{2}}{2\sigma_{a}^{2}})$. By the Bernstein inequality (Lemma B.9), we have
@@
Since $\|x\|_{2}^{2}=\sum_{i=1}^{d}x_{i}^{2}$, by Chi-square tail bound (Lemma B.7), we have
@@
Lemma E.3 (Sparsity analysis, formal version of Lemma 6.1).
@@
This proof follows from applying union bound on Lemma E.1 and Lemma E.2.
By Lemma E.2, we have
@@
Appendix F Running Time of Softmax Attention
@@
Theorem F.1 (Running time of Softmax attention generation, formal version of Theorem 4.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 1 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under the Softmax attention generation scenario: Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $m=\Theta(1)$. Algorithm 1 (replacing ReLU attention with Softmax attention) takes $O(mn^{4/5})$ time to generate the answer.
@@
The Softmax attention generation scenario can be proved by substituting the ReLU attention $\mathsf{Attn}_{r}$ (Definition 1.2) with Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ (Definition B.2) in Algorithm 1 and Theorem 4.1. ∎
@@
Theorem F.2 (Running time of Softmax full attention computation, formal version of Theorem 5.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 2 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under full Softmax attention computation scenario: Suppose we have $m=\Theta(n)$. Algorithm 2 (replacing ReLU attention with Softmax attention) takes $O(n^{2-1/\lfloor d/2\rfloor}d+n^{1+4/5}d)$ time to calculate the attention output.
@@
The Softmax full attention computation scenario can be proved by substituting the ReLU attention $\mathsf{Attn}_{r}$ (Definition 1.2) with Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ (Definition B.2) in Algorithm 2 and Theorem 5.1. ∎
Appendix G Error Analysis of Softmax Attention